# Personalized Generative Low-light Image Denoising and Enhancement

### v2

@@ Abstract
Modern cameras produce remarkably high-quality images, yet their performance in low-light conditions remains suboptimal due to fundamental limitations in photon shot noise and sensor read noise. While generative image restoration methods have shown promising results compared to traditional approaches, they often suffer from hallucinatory content generation when the signal-to-noise ratio (SNR) is low. Leveraging the availability of personalized photo galleries on users’ smartphones, we introduce Diffusion-based Personalized Generative Denoising (DiffPGD), a novel approach that builds a customized diffusion model for individual users. Our key innovation lies in the development of an identity-consistent physical buffer that extracts the physical attributes of the person from the gallery. This ID-consistent physical buffer serves as a robust prior that can be seamlessly integrated into the diffusion model to restore degraded images without the need for fine-tuning. Over a wide range of low-light testing scenarios, we show that DiffPGD achieves superior image denoising and enhancement performance compared to existing diffusion-based denoising approaches.
Figure 1: Using gallery from a phone user, we can restore low-light noisy facial images. Our method produces finer details and better identity consistency compared to other state-of-the-art restoration approaches like MIRNet [61, 62], GFP-GAN [52], DiffLL [22] and CodeFormer [67] on both simulated (top row) and real-captured (bottom row) low-light noisy inputs.
22footnotetext: Project page: https://genai-restore.github.io/DiffPGD/.
1 Introduction
With the astonishing development of smartphones, it is safe to say that smartphone cameras today have surpassed digital single-lens reflex (DSLR) cameras in both popularity and functionality. However, the small form factor of smartphone cameras puts a tight constraint on the aperture size, hence limiting the amount of light a CMOS pixel can detect. In a low-light imaging environment, this creates a fundamental limitation to how short the minimum exposure needs to be and how much signal-to-noise ratio (SNR) the sensor can support. While some of the noise seen at low light can be mitigated using better CMOS technology (e.g., correlated double sampling to reduce the read noise [41] and deep-well pump gate to reduce the stray capacitance between the floating diffusion and the transfer gate [39, 40]), the Poisson / Bose-Einstein statistics due to random photon arrivals is a problem created by mother nature that cannot be solved by even the ideal sensors. Therefore, image denoising and enhancement by exploiting the internal structures of images become a necessary task for all smartphone cameras.
@@ 1 Introduction
$\mathbf{y}=\text{ADC}\big{(}\text{Poisson}(\alpha\mathbf{x}+\eta)\big{)}+\epsilon,$ (1)
where $\mathbf{x}$ is the unknown clean image, $\alpha$ is the sensor gain, $\eta$ is the dark current, and $\epsilon$ is the read noise. The inverse problem associated with low-light denoising is to recover $\mathbf{x}$ from $\mathbf{y}$. Over the past five decades, the main driving force is to exploit image structures, from Wiener filter [36], total variation [44, 4], wavelets [9] to non-local means [3], BM3D [10], and recently convolution neural networks [25, 64, 50, 6, 2, 18, 57], transformers [35, 38, 63], and diffusion models [26, 60, 68]. A common theme across existing methods is to find and utilize a generic prior $p(\mathbf{x})$, i.e., a prior learned from a large collection of example images. While they perform well in moderately difficult problems, the generality of these methods cannot be extended to heavily corrupted images. Particularly in the context of human face recovery, the restored results often lack real identity and exhibit artifacts. This is attributed to the ill-posed nature of the problem and the lack of appropriate constraints.
Why Gallery Photos? Smartphone cameras today typically store hundreds if not thousands of a user’s photos, captured at different times, in different places, and under different lighting conditions. While these images have many variations, they are all about the same person(s). Therefore, if the imaging goal is to take a photo of this user, the gallery on the phone would be the best source to build a prior $p(\mathbf{x})$. This concept is illustrated in Fig. 2. In the context of diffusion-based image restoration, the initial solution space can be large since many candidate solutions are consistent with the noisy observation. The gallery provides a strong constraint to the search problem. This allows us to search for better quality images and preserve the person’s identity.
@@
  • How to incorporate the class-specific $p(\mathbf{x})$ into the restoration model: The bigger (and harder) question is that given the gallery photos, how do we efficiently extract the prior information and improve the restoration? We do not want to train a restoration model from the scratch because the gallery can be small. We want to avoid fine-tuning as much as possible during inference. Moreover, the person’s identity needs to be preserved.
Physical Buffers to the Rescue. Given the gallery photos, what kind of prior information would be useful for restoration? Advancements in computer vision have made it possible to extract facial physical buffers—including albedo and normal maps—from a person’s gallery of images[43, 65, 47]. These physical buffers capture crucial identity-defining properties such as surface geometry and skin color, effectively encoding an individual’s unique identity. At the same time, it also eliminates the external influence of environmental lighting, pose, and other identity-independent variables. By incorporating this robust prior that disentangles the external influences, our approach can effectively extracts identity information from the gallery.
Contributions. Our contributions are threefold:
@@
    Personalized Low-Light Denoise and Enhancement Framework: We introduce a new framework that leverages diffusion models and person-specific priors from photo galleries to denoise low-light facial images while presering the identity.
@@
    Physical Identity Buffers: We propose the use of physical identity buffers from existing face galleries to enrich the restoration process, enabling accurate reconstruction of human faces under severe degradation.
  3. 3.
    Easy Deployment: Our method operates easily without requiring fine-tuning for individual users, making it practical and scalable to multiple users.
Extensive experiments demonstrate that our approach outperforms existing techniques in both visual quality and identity preservation, providing a significant advancement in low-light facial image restoration.
2 Related Works
@@
Low-light Image Enhancement and Denoising. Low-light images often suffer from significant noise due to low signal-to-noise ratios (SNR), in addition to reduced brightness and contrast. Traditional low-light denoising techniques often focus on exploiting image structures, ranging from early methods like Wiener filtering [36], total variation [44, 4], and wavelets [9] to non-local means [3] and BM3D [10]. More recently, deep learning approaches, including convolutional neural networks [25, 64, 50, 6, 2, 18, 57, 66], transformers [35, 38, 63], and diffusion-based models [26, 60, 68, 42], have emerged as promising solutions for low-light denoising. Low-light image enhancement (LLIE) aims to improve the brightness and contrast of images captured in poorly lit conditions. Traditional approaches [16, 31, 29] often leverage heuristic or physical models to enhance image quality, while learning-based methods [28, 56, 24, 5, 17, 37, 53, 59, 22] have demonstrated greater effectiveness when supported by large-scale datasets. Recent advancements [45, 38, 42, 58] in this field have increasingly focused on jointly addressing denoising and enhancement tasks, aiming to improve image quality while effectively suppressing noise to maintain natural appearance. However, to the best of our knowledge, no specific approach currently exists for joint enhancement and denoising of facial images in low-light conditions.
Figure 3: Path to the idea. Using facial physical buffers and gallery images can help preserve human face identity information.
@@
Deep Face Restoration. CNN/Transformer-based face restoration methods [54, 14, 23], such as RestoreFormer [54], typically rely on paired low- and high-quality image datasets to learn resolution enhancement and noise reduction. While effective, these methods often struggle under severe degradation and lack the capability to synthesize realistic details in complex scenarios.
@@
In recent years, generative priors have been increasingly utilized in face restoration to better address real-world degradations. GAN-based methods, such as GFP-GAN [52], leverage StyleGAN2 [49] to enhance facial details, while diffusion-based models [55, 7, 13, 48, 51] have further expanded restoration capabilities. For instance, BFRDiffusion [7] uses Stable Diffusion to enhance low-quality images by adding high-fidelity details, and Ding et al. [13] employ a pretrained diffusion model that denoises degraded inputs while preserving identity by fine-tuning on a person’s gallery photos. MGBFR [48] combines text prompts and reference images with a dual-control adapter to retain accurate facial attributes and identity. Although these generative methods are effective, they often rely on multi-stage training or finetunning, which can limit the efficient use in some real-world applications.
@@
Reference Prior for Face Restoration. In face restoration, reference images [48, 13, 32, 33] provide essential priors for recovering detailed and high-frequency facial features. These priors guide the model to generate realistic restorations. Reference-based methods can be divided into two types: single reference and gallery-based images. The single reference approach [48] uses one image to guide the restoration, focusing on enhancing facial details from a specific reference. In contrast, the gallery-based approach [13] leverages multiple reference images, offering more diverse information and stronger constraints, improving the accuracy and realism of the restoration. Despite the effectiveness [13] using gallery photos of users, it still requires multi-stage training and fine-tuning on each user during inference.
3 Method
Figure 4: For images affected by mild low-light noise, phycial buffers can be accurately extracted. However, when encountering the images degraded by severe low light and noise, precise physical buffers are difficult to obtained.
3.1 Restoration by Diffusion
A diffusion model approximates the clean image distribution $p_{\theta}(\mathbf{x}_{0})$ by learning a model $\theta$ that reverses the noise-adding process. In Denoising Diffusion Probabilistic Models (DDPMs) [21], Gaussian noise is progressively added to a clean image $\mathbf{x}_{0}$:
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\mathbf{\epsilon},\quad\text{where}\quad\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (2)
@@ 3.1 Restoration by Diffusion
The reverse process then removes the noise to reconstruct the clean image. After training, for any given time $t$ and the corresponding noisy image $\mathbf{x}_{t}$, the model can iteratively denoise by sampling from $p(\mathbf{x}_{0}|\mathbf{x}_{t})$.
In restoration tasks, like low-light denoising, the goal is to recover a high-quality image $\mathbf{x}_{0}$ from a low-quality input $\mathbf{y}$. Unlike pure generation, where the output can be any image that follows the distribution of clean images, restoration is constrained by the content of $\mathbf{y}$. During training, the model uses $\mathbf{y}$ as a condition, and once the model is trained, for any given time $t$ and the corresponding noisy image $\mathbf{x}_{t}$, it can iteratively denoise by sampling from $p(\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y})$ using the trained model.
Figure 5: The overall architecture of the proposed method. Our core idea is to use ID-consistent physical buffers, extracted from gallery photos, to constrain the generative space in the diffusion model restoration process. For a high-quality gallery, we use LAP [65] to extract the albedo and normal information for each photo and apply adaptive aggregation to fuse the entire gallery. The extracted albedo represents base skin color and facial appearance, while normal captures facial geometry. In our framework, the output physical buffers isolate the intrinsic ID properties from lighting, shading, and pose, enabling the diffusion model to apply only ID-related information consistently.
3.2 Why Combining Gallery and Physical Buffer?
We formalize our core approach of extracting identity-consistent physical buffers from gallery photos to guide the diffusion model for restoration. This method is based on the observation that using facial physical buffers or gallery images can help preserve identity information during restoration tasks involving diffusion models. To evaluate the capability of physical buffers and gallery images, we conducted four simple experiments under a blind scheme (excluding $y_{0}$ during training), as illustrated in Fig. 3:

### v3

@@ Abstract
Modern cameras’ performance in low-light conditions remains suboptimal due to fundamental limitations in photon shot noise and sensor read noise. Generative image restoration methods have shown promising results compared to traditional approaches, but they suffer from hallucinatory content generation when the signal-to-noise ratio (SNR) is low. Leveraging the availability of personalized photo galleries of the users, we introduce Diffusion-based Personalized Generative Denoising (DiffPGD), a new approach that builds a customized diffusion model for individual users. Our key innovation lies in the development of an identity-consistent physical buffer that extracts the physical attributes of the person from the gallery. This ID-consistent physical buffer serves as a robust prior that can be seamlessly integrated into the diffusion model to restore degraded images without the need for fine-tuning. Over a wide range of low-light testing scenarios, we show that DiffPGD achieves superior image denoising and enhancement performance compared to existing diffusion-based denoising approaches. Our project page can be found at https://genai-restore.github.io/DiffPGD/.
I Introduction
With the astonishing development of smartphones, it is safe to say that smartphone cameras today have surpassed digital single-lens reflex (DSLR) cameras in both popularity and functionality. However, the small form factor of smartphone cameras puts a tight constraint on the aperture size, hence limiting the amount of light a CMOS pixel can detect. In a low-light imaging environment, this creates a fundamental limitation to how short the exposure can be and how much signal-to-noise ratio (SNR) the sensor can support. While some of the noise seen at low light can be mitigated using better CMOS technology (e.g., correlated double sampling to reduce the read noise [Nakamura2021sensor] and deep-well pump gate to reduce the stray capacitance [ma2015pump, Ma2022qis]), the Poisson statistics due to random photon arrivals is a problem created by mother nature that cannot be solved by even the ideal sensors. Therefore, image denoising and enhancement by exploiting the internal structures of images become a necessary task for all smartphone cameras.
@@ I Introduction
$\mathbf{y}=\text{ADC}\big(\text{Poisson}(\alpha\mathbf{x}+\eta)\big)+\text{Gauss}(0,\sigma_{\text{read}}^{2}),$ (1)
where $\mathbf{x}$ is the unknown clean image, $\alpha$ is the quantum efficiency, $\eta$ is the dark current, $\mathrm{ADC}(\cdot)$ models the analog-to-digital conversion that quantizes the sensor’s analog voltage into digital pixel values, and $\text{Gauss}(0,\sigma_{\text{read}}^{2})$ is the read noise. The inverse problem associated with low-light denoising is to recover $\mathbf{x}$ from $\mathbf{y}$. Over the past five decades, the main driving force is to exploit image structures, from Wiener filter [Lim1989TwoDimensionalSA], total variation [Rudin1994TV, Chambolle2004tv], wavelets [Coifman1995] to non-local means [Buades2005nl], BM3D [Dabov2007BM3D], and recently convolution neural networks [Jin2017cnn, Guo2019real, wei2020physics, chi2020dynamic, Elgendy2021LowLightDA], transformers [Liang2021swinir, Lu2022low, zhang2024xformer], and diffusion models [kawar2022denoising, zhu2023DiffPIR, Wang2025GenerativePB, Yuan2025iHDRIH]. A common theme across existing methods is to find and utilize a generic prior $p(\mathbf{x})$, i.e., a prior learned from a large collection of example images. While they perform well in moderately difficult problems, the generality of these methods cannot be extended to heavily corrupted images. In the context of human face recovery, the restored results often lack real identity and exhibit artifacts. This is attributed to the ill-posed nature of the problem and the lack of appropriate constraints.
Why Gallery Photos? Smartphone cameras today typically store hundreds if not thousands of a user’s photos, captured at different times, in different places, and under different lighting conditions. While these images have many variations, they are all about the same person(s). Therefore, if the imaging goal is to take a photo of this user, the gallery on the phone would be the best source to build a prior $p(\mathbf{x})$. This concept is illustrated in Fig. 1. In this paper, we are interested in diffusion-based image restoration. The role of gallery is to provide a strong constraint to the search problem. This allows us to search for better quality images and preserve the person’s identity.
@@
  • How to incorporate the class-specific $p(\mathbf{x})$ into the restoration model: This is a bigger (and harder) question. Given the gallery photos, how do we efficiently extract the prior information and improve the restoration? We do not want to train a restoration model from scratch because the gallery can be small. We also want to avoid fine-tuning during inference. Moreover, the person’s identity needs to be preserved.
Physical Buffers to the Rescue. Given the gallery photos, what kind of prior information would be useful for restoration? Advancements in computer vision have made it possible to extract facial physical buffers—including albedo and normal maps—from a person’s gallery of images[roth2016adaptive, Zhang2021lap, suwajanakorn2015makes]. These physical buffers capture crucial identity-defining properties such as surface geometry and skin color, effectively encoding an individual’s unique identity. At the same time, it also eliminates the external influence of environmental lighting, pose, and other identity-independent variables. By incorporating this robust prior that disentangles the external influences, our approach can effectively extract identity information from the gallery.
Contributions. Our contributions are as follows:
@@
    We introduce a personalized low-light denoising framework based on diffusion models. The proposed framework extracts person-specific priors from gallery images and integrates them into the diffusion process to effectively restore low-light facial images. Our solution offers three advantages: (i) We outperform existing methods in low-light conditions. (ii) We preserve person identity better than existing methods. (iii) We do not need to fine-tune for individuals.
@@
    We propose to leverage physical buffer as an explicit prior for identity-preserving restoration. By learning the physical prior from clean data rather than noisy observations, the model captures intrinsic and noise-independent identity features. This physically grounded prior demonstrates strong generalization capability—models trained on synthetic data can effectively transfer to real-world scenarios.
II Related Works
@@
Low-light Image Enhancement and Denoising. Low-light images often suffer from significant noise due to low signal-to-noise ratios (SNR), in addition to reduced brightness and contrast. Traditional low-light denoising techniques often focus on exploiting image structures, ranging from early methods like Wiener filtering [Lim1989TwoDimensionalSA], total variation [Rudin1994TV, Chambolle2004tv], and wavelets [Coifman1995] to non-local means [Buades2005nl] and BM3D [Dabov2007BM3D]. More recently, deep learning approaches, including convolutional neural networks [Jin2017cnn, Guo2019real, wei2020physics, zheng2021adaptive, Fang2025SRENetSL, chi2020dynamic, Elgendy2021LowLightDA], transformers [Liang2021swinir, Lu2022low, zhang2024xformer, Yang2025LearningTS], and diffusion-based models [kawar2022denoising, zhu2023DiffPIR, Nguyen2024diffusion, CheukYiuChan2024AnlightenDiffAD, Wu2025MutuallyRL, Prateek2025quantadiff], have emerged as promising solutions for low-light denoising. Low-light image enhancement (LLIE) aims to improve the brightness and contrast of images captured in poorly lit conditions. Traditional approaches [Fu2016Weighted, Li2018Structure, Lee2013Contrast] often leverage heuristic or physical models to enhance image quality, while learning-based methods [Land1971Retinex, Chen2018Retinex, Jiang2021EnlightenGAN, Chen2018See, Guo2020Zero, Lu2021Tbefn, wang2023exposurediffusion, xu2022snr, jiang2023low] have demonstrated greater effectiveness when supported by large-scale datasets. Some works also adopt training-free (zero-shot) learning for LLIE [Lv2024FourierPD, Huang2025ZeroShotLI, Xie2024ResidualQL]. Recent advancements [Shi2024ZERO, Lu2022low, Nguyen2024diffusion, xu2020learning] in this field have increasingly focused on jointly addressing denoising and enhancement tasks, aiming to improve image quality while effectively suppressing noise to maintain natural appearance. However, to the best of our knowledge, no specific approach currently exists for joint enhancement and denoising of facial images in low-light conditions.
@@
Deep Face Restoration. CNN/Transformer-based face restoration methods [wang2022RestoreFormer, Fan2022faceformer, Jiang2022dual], such as RestoreFormer [wang2022RestoreFormer], typically rely on paired low- and high-quality image datasets to learn resolution enhancement and noise reduction. CodeFormer [zhou2022towards] formulates face restoration as a code prediction task in a discrete latent space, leveraging a Transformer to recover identity-consistent and perceptually realistic results. While effective, these methods often struggle under severe degradation and lack the capability to synthesize realistic details in complex scenarios.
@@
In recent years, generative priors have been increasingly utilized in face restoration to better address real-world degradations. GAN-based methods, such as GFP-GAN [wang2021towards], leverage StyleGAN2 [tov2021stylegan2] to enhance facial details, while diffusion-based models [wang2023dr2, Chen2024face, ding2024restoration, Tao2024face, Varanka2024gallery] have further expanded restoration capabilities. For instance, BFRDiffusion [Chen2024face] uses Stable Diffusion to enhance low-quality images by adding high-fidelity details, and Ding et al. [ding2024restoration] employ a pretrained diffusion model that denoises degraded inputs while preserving identity by fine-tuning on a person’s gallery photos. MGBFR [Tao2024face] combines text prompts and reference images with a dual-control adapter to retain accurate facial attributes and identity. Although these generative methods are effective, they often rely on multi-stage training or fine-tuning and often struggle recovering identity-consistent face under severe degradation, which can limit the efficient use in some real-world applications.
@@
Reference Prior for Face Restoration. In face restoration, reference images [Tao2024face, ding2024restoration, li2022learning] provide essential priors for recovering detailed and high-frequency facial features. These priors guide the model to generate realistic restorations. Reference-based methods can be divided into two types: single reference and gallery-based images. The single reference approach [Tao2024face] uses one image to guide the restoration, focusing on enhancing facial details from a specific reference. In contrast, the gallery-based approach [ding2024restoration] leverages multiple reference images, offering more diverse information and stronger constraints, improving the accuracy and realism of the restoration. Despite the effectiveness [ding2024restoration] using gallery photos of users, it still requires multi-stage training and fine-tuning on each user during inference.
The summary comparison between our method and previous methods is summarized in Table I, and the preview of the visual results can be found in Fig. 2.
III Method
Our Proposed method is a generative approach based on diffusion models (Section III-A). However, a naive diffusion model does not support a personalized gallery as prior. The core innovation of our proposed method is the introduction of a physical buffer that can be extracted from the personalized gallery (Section III-B and Section III-C), and then integrated into the diffusion models to achieve the target denoising (Section III-D).
III-A Denoising Diffusion Model
Diffusion models [SohlDickstein2015DeepUL, ho2020ddpm] approximate a complex data distribution $q(\mathbf{x}_{0})$ by learning to reverse a gradual noising process that transforms the data into a Gaussian distribution $\mathcal{N}(0,\mathbf{I})$. Through this iterative denoising process, they have become the state of the art in generative modeling, outperforming GANs and VAEs in stability and visual fidelity.
Forward diffusion. Given a clean image $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$, the forward process progressively corrupts it with Gaussian noise of increasing variance $\{\beta_{t}\}_{t=1}^{T}$, forming a Markov chain:
$q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}\!\big(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\,\beta_{t}\mathbf{I}\big).$ (2)
After $T$ steps, the image becomes nearly indistinguishable from Gaussian noise. Using properties of Gaussian composition, the sample at an arbitrary step can be expressed as
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,\mathbf{I}),$ (3)
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ controls the remaining signal-to-noise ratio. When $\bar{\alpha}_{t}$ approaches zero, the original signal is nearly destroyed. Typical choices of the variance schedule include the linear schedule of Ho et al. [ho2020ddpm, dhariwal2021diffusion] and the cosine schedule of Nichol et al. [Nichol2021ImprovedDD], which better distribute the noise levels across timesteps and improve sample quality.
TABLE I: Compared to previous related methods, our approach performs joint low-light enhancement and denoising (low-light), effectively recovering identity-consistent faces even under severe degradation (ID-consistent). By leveraging the user’s gallery, our method supports personalized restoration (Personalized), and is scalable to multiple users without requiring fine-tuning when switching users (No fine-tune).
Method low-light ID-consistent Personalized No fine-tune
DiffLL [jiang2023low]
FourierDiff [Lv2024FourierPD]
MIRNet [zamir2020learning]
SNR-aware [xu2022snr]
GFPGAN [wang2021towards]
CodeFormer [zhou2022towards]
Ding et al. [ding2024restoration]
PFStorer [Varanka2024gallery]
DiffPGD (Ours)
Reverse denoising. The generative (reverse) process learns to invert the corruption by gradually removing noise from $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$ to recover $\mathbf{x}_{0}$. A neural network $\epsilon_{\theta}$ is trained to predict the noise component added at each step, thereby parameterizing the reverse conditional
$p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})=\mathcal{N}\!\big(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t,\mathbf{c}),\sigma_{t}^{2}\mathbf{I}\big),$ (4)
where $\mathbf{c}$ denotes any optional conditioning signal. The network is optimized via the simplified denoising objective [ho2020ddpm, rombach2022high] derived from the variational lower bound:
$\mathcal{L}=\mathbb{E}_{\mathbf{x}_{0},\,\mathbf{c},\,\epsilon,\,t}\big[\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})\|_{2}^{2}\big],$ (5)
where $t$ is uniformly sampled from $\{1,\ldots,T\}$. Intuitively, $\epsilon_{\theta}$ learns a vector field that points from noisy samples back to the data manifold at different noise scales.
@@ III-A Denoising Diffusion Model
Sampling process. At inference, given a noisy image $\mathbf{x}_{t}$, the model iteratively refines it by sampling from the learned reverse process:
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{c})\right)+\sigma_{t}z,$ (6)
where $z\sim\mathcal{N}(0,\mathbf{I})$ and typically $\sigma_{t}^{2}=\beta_{t}$. This iterative denoising gradually reconstructs the clean data $\mathbf{x}_{0}$ from pure noise $\mathbf{x}_{T}$.
This forward–reverse framework provides a simple yet powerful formulation that unifies generative modeling, denoising, and Bayesian inference. By adjusting the conditioning signal $\mathbf{c}$, the same architecture can handle diverse tasks such as image restoration, super-resolution, and text-guided generation. In restoration settings, $\mathbf{c}$ often represents the degraded observation or auxiliary features (e.g., edge, illumination, or identity), and the reverse process recovers a clean image consistent with this condition.
Figure 2: Preview of the visual results. Using gallery photos from a user’s smartphone, we can restore low-light, noisy facial images. Our method produces finer details and better identity compared to state-of-the-art low-light and face restoration approaches like MIRNet [zamir2020learning], DiffLL [jiang2023low], GFP-GAN [wang2021towards], FourierDiff [Lv2024FourierPD], and CodeFormer [zhou2022towards]. Please zoom in for the best visual experience.
III-B Why Combining Gallery and Physical Buffer?
Motivation. After introducing the diffusion framework in Section III-A, we recall that a conditional diffusion model requires additional information $\mathbf{c}$ to constrain the reverse process. For personalized restoration, such a condition must encode the subject’s identity to prevent hallucination. We therefore investigate two possible priors for identity preservation: (1) the gallery prior, and (2) the physical-buffer prior.
A gallery refers to a small set (typically 3–6) of high-quality face photos of the target person, captured under different poses and lighting conditions. These images are readily available from user photo collections. They contain rich identity cues, but are also affected by background, expression, and illumination variations.
A physical buffer refers to an intermediate facial representation—in our case the albedo and surface normal—that describes the physical properties of a person’s face independent of lighting and shading. Albedo captures the intrinsic color of an object’s surface—that is, the subject’s color properties without any lighting effects. It reveals a person’s facial appearance, e.g. eyebrows, lip coloration, and other distinct texture elements can reflect individual identity traits. Surface normal encodes the surface orientation at each pixel, which provides hints about the underlying shape characteristics. These buffers can be extracted using pretrained face reconstruction networks [feng2021learning, Zhang2021lap]. Compared to the gallery, they provide disentangled, lighting-independent identity features, but are less directly visual.
Controlled Comparison. To analyze the complementary effects of these priors, we conduct four controlled generative experiments (Fig. 3).